# Leveraging Large Language Models to Measure Gender Bias in Gendered Languages

### v1

@@ 4.1 Datasets
We evaluate the proposed method on several parallel datasets from the OPUS Machine Translation project dataset collection [^1] The website specifies that the data provided on the platform are believed to be free to redistribute. (Tiedemann, 2012) that includes aligned translations between Spanish and English. We focus on parallel datasets to facilitate the comparison of the gender polarity method for English with our LLM-based gender bias quantification method for Spanish. The datasets were chosen based on their extended use in machine translation and NLP model training in general. In particular, we evaluated our approach on the Europarl, CCAligned, Global Voices, and WMT-News datasets.
@@ 4.2 Validation
We evaluated the following LLMs available through the OpenAI API: gpt-4-turbo-preview (gpt-4-0125-preview), gpt-4o (gpt-4o-2024-05-13), and gpt-4-turbo (gpt-4-turbo-2024-04-09). For each of them, we executed five runs of our method with the same (default) configuration. For few-shot prompting, we used five sentence-analysis pairs listed in Appendix B. The budget spent on these experiments was approximately 10 USD.
To evaluate the correctness of the LLM output, we employed a case-insensitive comparison of the identified words and the (mis)match of the two attributes ($p$ and $g$, see Section 3.3) w.r.t. the ground truth. We report the number of words that were correctly identified and correctly classified in both attributes ($n_{c}$), correctly identified but incorrectly classified in at least one attribute ($n_{i}$), missed (not identified) by the method ($n_{m}$), and extra words that do not appear in the ground truth but were returned by the method ($n_{e}$). Using these values, we define the following performance metrics:
@@
As reflected in Table 1, gpt-4-turbo performed the best across all metrics. Hence, we selected gpt-4-turbo as the LLM of choice for our empirical study. In addition, we tested four open-source models from the Ollama [^2] https://github.com/ollama/ollama project that can be run for inference locally: Llama 2 Uncensored (7B), Mistral (7B), Neural Chat (7B), and Llama 2 (13B). Unfortunately, all of them were found unsuitable due to their poor performance. This task appears to be too complex for these smaller models, as they did not even output the analysis in the specified format in the vast majority of cases.
@@ 4.3 Results
Table 2: Evaluation of two representative samples for each of the four benchmark datasets in English using the gender polarity method. The reported results follow the notation introduced in Section 3.1. The gender bias (male:female) is reported in the last column.
@@
Table 3: Evaluation of two representative samples for each of the four benchmark datasets in Spanish using the proposed LLM-based method. The reported results follow the notation introduced in Section 3.3, and the asterisk denotes any value of the binary variable. The gender bias (male:female) is reported in the last column.
@@
While the male:female ratio varies significantly across datasets, all datasets are biased towards masculine references in both languages. The overall agreement of the results for the two subsets of each dataset suggests that the subsets are well-representative. The results indicate that the gender bias is substantially more pronounced in Spanish than in English, although this comparison needs to be treated with caution as the methods are different.
Interestingly, in Spanish, the lowest male:female disparity is found in the Europarl dataset, which makes intuitive sense as the European Parliament is a political institution that strives for gender equality and representation across its member states, reflecting broader societal efforts towards balanced gender representation in government and public discourse. Conversely, the largest male:female disparity is identified in the WMT-News dataset, suggesting that news reporting and journalism – especially in certain regions or topics – may still be male-dominated, reflecting broader societal trends and biases in media representation and participation.
These insights highlight the importance of using diverse data sources to train and evaluate NLP models to ensure that they do not perpetuate existing biases. The results confirm the utility of our LLM-based method in capturing and quantifying gender biases effectively.
@@ 5 Discussion
The results of our study have significant implications for the field of natural language processing, particularly in the understanding and mitigation of gender bias in gendered languages. By employing state-of-the-art LLMs to analyze Spanish corpora, our research highlights the extent to which gender bias can permeate linguistic datasets and subsequently influence machine learning models trained on these datasets. Below, we summarize the main findings of our research.
1. LLMs are a promising tool to identify gender bias in gendered corpora
There is a lack of methods to automatically detect gender disparity in text corpora in gendered languages, such as Spanish. To the best of our knowledge, we are the first to propose and empirically demonstrate the value of LLMs, especially the larger and more advanced versions like gpt-4-turbo, as a tool to effectively identify and classify gendered language within a complex linguistic framework like Spanish. This capability is crucial for developing fairer NLP applications, as it allows for a deeper understanding of how gender is represented in text and how these representations can lead to biased outcomes.
2. Gender bias pervades commonly used corpora in Spanish
We identify a significant male:female gender disparity, ranging between 4:1 (Europarl dataset) and 5–6:1 (WMT-News dataset), in commonly used benchmark datasets in Spanish. The Europarl dataset exhibits the smallest gender disparity, which is aligned with the institution’s commitment to gender equality and balanced representation. This finding suggests that legislative efforts and policies aided at promoting gender balance can significantly influence the language and content of official records. However, the number of references to males is four times the number of references to females, indicating that more efforts are needed to achieve representative gender equality. Conversely, we identify the largest gender disparity in the WMT-News dataset, which suggests that news reporting and journalism remain male-driven, reflecting broader societal trends and editorial biases in the media industry.
The prevalence of masculine w.r.t. feminine references across different datasets underscores the pervasive nature of gender bias in text corpora and the importance of identifying and mitigating such biases to create more balanced datasets, which in turn would enable more equitable models.
@@ 2. Gender bias pervades commonly used corpora in Spanish
The differences in gender ratio between the English and Spanish versions of our datasets further underline the linguistic and cultural variances affecting gender representation. While both languages exhibited biases, the Spanish datasets generally showed a higher degree of male dominance. This may be partly due to the grammatical structure of Spanish, which includes gendered nouns and more explicit gender marking, potentially amplifying the visibility and perpetuation of gender biases. The English datasets, while also biased, tended to have lower ratios of male:female references, possibly due to a combination of linguistic factors and different societal norms influencing the text. Interestingly, in the English corpora, the CCAligned samples exhibit the lowest gender disparity. Conversely and aligned with the findings in the Spanish case, the WMT-News samples have the largest male:female ratio. These results highlight the need for considering both linguistic and cultural contexts when analyzing and mitigating biases in multilingual and multicultural corpora.
@@ 6 Conclusion
We have presented a novel methodology for measuring gender bias in Spanish text corpora using large language models. Through experiments with four benchmark datasets, we have demonstrated the efficacy of this approach, which leverages the capabilities of LLMs to differentiate between gendered references in a nuanced manner.
Our findings reveal significant gender disparities that range between 4:1 and 6:1 in the male-to-female ratio, depending on the dataset, highlighting the need for continuous monitoring and correction of biases in linguistic data sources. The methodology developed herein not only contributes to the field of bias research in NLP but also sets a precedent for future studies aiming to understand and rectify gender bias in languages other than English, broadening the scope of bias detection and mitigation strategies to a global context. Ensuring that language technologies do not perpetuate or even amplify existing social biases is essential.
In future work, we aim to further refine our approach and explore the integration of additional linguistic and cultural nuances that may affect the representation of gender in text. Moreover, extending this framework to other gendered languages could provide a more comprehensive global perspective on gender bias in NLP. Continued advancements in model capabilities and a growing understanding of bias will undoubtedly enhance the fairness and accuracy of language-based AI systems.
@@ Acknowledgments
The work of E.D., S.S., and N.O. on this paper was supported by a nominal grant received at the ELLIS Unit Alicante Foundation from the Regional Government of Valencia in Spain (Convenio Singular signed with Generalitat Valenciana, Conselleria de Innovación, Industria, Comercio y Turismo, Dirección General de Innovación). E.D. is also supported by Intel Corporation.
The work of Y.G. and P.M. on this paper was partially funded by the University of Alicante, the Spanish Ministry of Science and Innovation, the Generalitat Valenciana, and the European Regional Development Fund (ERDF) through the following funding: As main, VIVES: “Pla de Tecnologies de la Llengua per al valencià” project (2022/TL22/00215334) from the Projecte Estratègic per a la Recuperació i Transformació Econòmica (PERTE) and the following supportive projects: NL4DISMIS (CIPROM/2021/21) with funds granted by the Generalitat Valenciana (Conselleria d’Educacio, Investigacio, Cultura i Esport), COOLANG (PID2021-122263OB-C22), and CORTEX (PID2021-123956OB-I00) with funds granted by MCIN/AEI/10.13039/501100011033 and, as appropriate, by “ERDF A way of making Europe”, by the “European Union” or by the “European Union NextGenerationEU/PRTR”.
@@
While we believe that our study provides valuable insights into the measurement of gender bias in gendered languages using LLMs, we have identified several limitations.
Inclusion of epicene words: In our analysis, we included epicene words – nouns that have the same form, including grammatical gender, for male and for female referents. For example, ‘la persona’ (person) is grammatically feminine but can refer to individuals of any gender. This inherent ambiguity can lead to inaccuracies in gender classification. In the data analyzed in this work, epicene words represent 5.8 % of all identified person-referencing words. It is noteworthy that 73.7 % of epicene words were counted as feminine (see Appendix C), such that excluding the epicene words would make the gender ratios even more biased towards the masculine gender.
Contextuality of the analysis: Our methodology focuses primarily on the occurrence and classification of gendered nouns and pronouns without deeply analyzing the context in which these words appear. This approach provides a quantitative measure of gender representation but lacks the nuance of qualitative analysis. Understanding the context could reveal more about the nature of gender bias, such as whether certain roles or characteristics are predominantly associated with a specific gender. Future work should incorporate contextual analysis to provide a richer and more comprehensive understanding of gender bias in texts.
@@
Budget implications: Using commercial, state-of-the-art LLMs in research can have a non-negligible financial cost. While we selected representative subsets from widely used benchmark datasets, a more extensive analysis could potentially provide more robust and generalizable findings at a cost.
@@
Ethics Statement
We aim to promote fairness and inclusivity by identifying and quantifying biases in text corpora used to train LLMs. We have adhered to ethical standards and code of conduct by ensuring transparency, reproducibility, and validating our methodology against manually annotated data. The datasets used for evaluation are publicly available benchmark datasets. While our work highlights significant gender disparities, we recognize the limitations of focusing on grammar-based gender classification and the reliance on specific LLMs.
We are committed to ethical AI use, advocating for continuous improvement in bias detection and mitigation techniques. Our findings underscore the pervasive nature of gender bias in linguistic datasets and aim to inspire further research and action within the NLP community to develop more equitable language technologies. By raising awareness and providing robust methodologies, we contribute to the broader discourse on AI ethics and the pursuit of technologies that uphold human dignity and equality.

### v2

@@ 4.1 Datasets
We evaluate the proposed method on several parallel datasets from the OPUS Machine Translation project dataset collection [^2] The website specifies that the data provided on the platform are believed to be free to redistribute. (Tiedemann, 2012) that includes aligned translations between Spanish and English. We focus on parallel datasets to facilitate the comparison of the gender polarity method for English with our LLM-based gender representation bias quantification method for Spanish. The datasets were chosen based on their extended use in machine translation and NLP model training in general. In particular, we evaluated our approach on the Europarl, CCAligned, Global Voices, and WMT-News datasets.
@@ 4.2 Validation
We evaluated the following LLMs available through the OpenAI API: gpt-4-turbo-preview (gpt-4-0125-preview), gpt-4o (gpt-4o-2024-05-13), and gpt-4-turbo (gpt-4-turbo-2024-04-09). For each of them, we executed five runs of our method with the same (default) configuration. For few-shot prompting, we used five sentence-analysis pairs listed in Appendix B. The budget spent on these experiments was approximately 10 USD. Note that the objective of our work is not to compare the performance of different LLMs in detecting gender imbalances but rather to show that, with careful prompt engineering, state-of-the-art LLMs can accurately identify gender representation imbalances in corpora of gendered languages.
To evaluate the correctness of the LLM output, we employed a case-insensitive comparison of the identified words and the (mis)match of the two attributes ($p$ and $g$, see Section 3.3) w.r.t. the ground truth. We compute the number of words that were correctly identified and correctly classified in both attributes ($n_{c}$), correctly identified but incorrectly classified in at least one attribute ($n_{i}$), missed (not identified) by the method ($n_{m}$), and extra words that do not appear in the ground truth but were returned by the method ($n_{e}$). Using these values, we define the following performance metrics:
@@
As reflected in Table 1, gpt-4-turbo performed the best across all metrics. Hence, we selected gpt-4-turbo as the LLM of choice for our empirical study. In addition, we tested four open-source models from the Ollama [^3] https://github.com/ollama/ollama project that can be run for inference locally: Llama 2 Uncensored (7B), Mistral (7B), Neural Chat (7B), and Llama 2 (13B). Unfortunately, all of them were found unsuitable for this task due to their poor performance. This task appears to be too complex for these smaller models, as they did not even produce a meaningful output in the vast majority of cases.
@@ 4.3 Results
Table 2: Gender representation results on two representative samples for each of the four benchmark datasets in English using the gender polarity method. The reported results follow the notation introduced in Section 3.1. The gender representation bias (male:female) is reported in the last column.
@@
Table 3: Gender representation results on two representative samples for each of the four benchmark datasets in Spanish using the proposed LLM-based method. The reported results follow the notation introduced in Section 3.3, and the asterisk denotes any value of the binary variable. The gender representation bias (male:female) is reported in the last column.
@@
While the male:female ratio varies significantly across datasets, all datasets are biased towards masculine references in both languages. The overall agreement of the results for the two subsets of each dataset suggests that the subsets are well-representative. The results indicate that the gender representation bias is substantially more pronounced in Spanish than in English, although this comparison needs to be treated with caution as the methods are different.
These insights highlight the importance of using diverse data sources to train and evaluate NLP models to ensure that they do not perpetuate existing biases. The results confirm the utility of our LLM-based method in capturing and quantifying gender representation bias effectively.
@@ 5 Discussion
The results of our study have significant implications for the field of natural language processing, particularly in the understanding and mitigation of gender bias in gendered languages. By employing state-of-the-art LLMs to analyze Spanish corpora, our research highlights the extent to which gender representation bias can permeate linguistic datasets and subsequently influence machine learning models trained on these datasets. Below, we provide a discussion of the main findings of our research.
1. LLMs are a promising tool to identify gender representation bias in gendered corpora
There is a lack of methods to automatically detect gender representation bias in text corpora in gendered languages, such as Spanish. To the best of our knowledge, we are the first to propose and empirically demonstrate the value of LLMs, especially the larger and more advanced versions like gpt-4-turbo, as a tool to effectively identify and classify gendered language within a complex linguistic framework like Spanish. This capability is crucial for developing fairer NLP applications, as it allows for a deeper understanding of how gender is represented in text and how these representations can lead to biased outcomes.
2. Gender representation bias pervades commonly used corpora in Spanish
We identify a significant male:female gender representation disparity, ranging between 4:1 (Europarl dataset) and 5–6:1 (WMT-News dataset), in commonly used benchmark datasets in Spanish. The Europarl dataset exhibits the smallest gender representation disparity, which is aligned with the institution’s commitment to gender equality and balanced representation. This finding suggests that legislative efforts and policies aided at promoting gender balance can significantly influence the language and content of official records. However, the number of references to males is four times the number of references to females, indicating that more efforts are needed to achieve representative gender equality. Conversely, we identify the largest gender representation disparity in the WMT-News dataset, which suggests that news reporting and journalism remain male-driven, reflecting broader societal trends and editorial biases in the media industry.
The prevalence of masculine w.r.t. feminine references across different datasets underscores the pervasive nature of gender representation bias in text corpora and the importance of identifying and mitigating such biases to create more balanced datasets, which in turn would enable more equitable models.
@@ 2. Gender representation bias pervades commonly used corpora in Spanish
The differences in gender ratio between the English and Spanish versions of our datasets further underline the linguistic and cultural variances affecting gender representation. While both languages exhibited biases, the Spanish datasets generally showed a higher degree of male dominance. This may be partly due to the grammatical structure of Spanish, which includes gendered nouns and more explicit gender marking, potentially amplifying the visibility and perpetuation of gender biases. The English datasets, while also biased, tend to have lower ratios of male:female references, possibly due to a combination of linguistic factors and different societal norms influencing the text. Interestingly, in the English corpora, the CCAligned samples exhibit the lowest gender representation disparity. Conversely and aligned with the findings for Spanish, the WMT-News samples have the largest male:female ratio. These results highlight the need for considering both linguistic and cultural contexts when analyzing and mitigating biases in multilingual and multicultural corpora.
4. Representation bias and its impact
While considering the imbalance of gendered words in training corpora as a measure of bias can be viewed as simplifying the problem, it represents an essential first step, particularly in gendered languages, as it provides a quantitative assessment of representation bias. This type of bias is one of the three identified forms of gender bias in language, which include: ‘language that ignores women; language that defines women narrowly; and language that depreciates women’ (Martin, 2020; Weatherall et al., 2002). Our focus is on the first form – language that ignores women – by measuring the differences in the prevalence of words that refer to male versus female human entities. Discourse analysis has demonstrated that imbalanced portrayals or frequency of gender mentions can lead to representational biases, as language often reflects societal norms and values. An overrepresentation of male-oriented language implies a bias toward male perspectives, effectively underrepresenting women and making female perspectives less visible (Pauwels, 2003). By quantifying these imbalances, we can lay the foundation for more comprehensive bias-mitigation strategies, which can later be applied to address subtler forms of bias in model outputs and behavior (Zhao et al., 2017; Babaeianjelodar et al., 2020). Importantly, the main contribution of this paper is to provide a method to measure gender representation imbalances in gendered languages, filling a gap in existing research where such a tool has yet to be developed.
Our work is framed within a larger initiative aimed at training large language models from scratch in non-English languages, where it is crucial to assess and mitigate gender biases within the training corpora (Wolfe and Caliskan, 2021). Biases originating in training data, particularly an imbalance in male/female words, can lead to biased models that replicate and potentially amplify these biases. When training data is predominantly male-oriented, models are more likely to generate outputs that favor male markers and perspectives. As such, addressing these biases early in the model development process, as we propose, is critical to creating balanced and equitable language models for non-English languages.
5. Implications for debiasing strategies
While the primary contribution of our paper is a novel method for quantifying gender representation bias in corpora, it is important to recognize that accurate detection and quantification of bias is a foundational first step in the debiasing process. Without reliable methods to identify and quantify bias, subsequent debiasing efforts may lack precision and effectiveness. Our method fills a critical gap by offering a systematic and robust approach to detect biases that are often overlooked, particularly in gendered languages.
The insights gained from our methodology can directly inform and enhance debiasing strategies by identifying specific instances and patterns of bias in training corpora. For example, if our tool detects that a dataset consistently uses masculine references, such as ‘el doctor’ for ‘the doctor’, this insight enables the development of targeted interventions to mitigate the bias. Potential debiasing strategies that our method could support include:
  • Gender-Neutral Rewriting: Once gender-marked words and phrases are identified, they can be rewritten in a gender-neutral manner where contextually appropriate. For instance, ‘el doctor’ could be replaced with ‘el personal médico’ (the medical personnel), which avoids gender marking.
  • Balanced Representation: To balance gender representation, sentences with the opposite gender marker can be added. For example, if ‘el doctor’ predominates, additional sentences featuring the female version of the word, ‘la doctora’, could be incorporated to balance the dataset.
  • Fair Data Sampling: During training, the sampling probability of gendered terms can be dynamically adjusted to maintain gender balance. For instance, a lower sampling rate might be applied to more frequently occurring masculine terms, which helps prevent their overrepresentation in the trained model.
@@ 6 Conclusion
We have presented a novel methodology for measuring gender representation bias in Spanish text corpora using large language models. Through experiments with four benchmark datasets, we have demonstrated the efficacy of this approach, which leverages the capabilities of LLMs to differentiate between gendered references in a nuanced manner.
Our findings reveal significant gender representation disparities that range between 4:1 and 6:1 in the male-to-female ratio, depending on the dataset, highlighting the need for continuous monitoring and correction of biases in linguistic data sources. The methodology developed herein not only contributes to the field of bias research in NLP but also sets a precedent for future studies aiming to understand and rectify gender bias in languages other than English, broadening the scope of bias detection and mitigation strategies to a global context. Ensuring that language technologies do not perpetuate or even amplify existing social biases is essential.
While our current method focuses specifically on detecting structural biases, it represents a foundational step in a larger framework aimed at addressing various contextual bias types. This focus aligns with the ‘methodological pyramid’ described at the end of Section 2, where gendered markers, though not the only type of bias, are a well-documented and impactful source of gender bias in language models. In future work, we aim to expand upon this foundation to incorporate additional forms of contextual bias. Leveraging the semantic capabilities of LLMs, we plan to refine our approach to detect more subtle and complex gender bias patterns embedded in language. The insights gained from our focus on structural bias can guide broader strategies for bias quantification and mitigation, supporting the development of models that capture linguistic and social contexts with greater nuance. This progression will ultimately enable a more comprehensive approach to fair and ethical language model development.
@@ Acknowledgments
The work of E.D., S.S., and N.O. on this paper was supported by a nominal grant received at the ELLIS Unit Alicante Foundation from the Regional Government of Valencia in Spain (Convenio Singular signed with Generalitat Valenciana, Conselleria de Innovación, Industria, Comercio y Turismo, Dirección General de Innovación). E.D. and N.O. are also supported by Intel Corporation. E.D. is also supported by the VIVES: ‘Pla de Tecnologies de la Llengua per al valencià’ project (2022/TL22/00215334) from the Projecte Estratègic per a la Recuperació i Transformació Econòmica (PERTE).
The work of Y.G. and P.M. on this paper was partially funded by the University of Alicante, the Spanish Ministry of Science and Innovation, the Generalitat Valenciana, and the European Regional Development Fund (ERDF) through the following funding: As main, VIVES: ‘Pla de Tecnologies de la Llengua per al valencià’ project (2022/TL22/00215334) from the Projecte Estratègic per a la Recuperació i Transformació Econòmica (PERTE) and the following supportive projects: NL4DISMIS (CIPROM/2021/21) with funds granted by the Generalitat Valenciana (Conselleria d’Educacio, Investigacio, Cultura i Esport), COOLANG (PID2021-122263OB-C22), and CORTEX (PID2021-123956OB-I00) with funds granted by MCIN/AEI/10.13039/501100011033 and, as appropriate, by ‘ERDF A way of making Europe’, by the ‘European Union’ or by the ‘European Union NextGenerationEU/PRTR’.
@@
While we believe that our study provides valuable insights into the measurement of gender representation bias in gendered languages using LLMs, we have identified several limitations.
Inclusion of epicene words: In our analysis, we included epicene words – nouns that have the same form, including grammatical gender, for male and for female referents. For example, ‘la persona’ (person) is grammatically feminine but can refer to individuals of any gender. This inherent ambiguity can lead to inaccuracies in gender classification. In the data analyzed in this work, epicene words represent only 5.8 % of all identified person-referencing words. It is noteworthy that 73.7 % of epicene words were counted as feminine (see Appendix C), such that excluding the epicene words would make the gender ratios even more biased towards the masculine gender.
The role of context: Our methodology focuses primarily on the occurrence and classification of gendered nouns and pronouns without deeply analyzing the context in which these words appear. This approach provides a quantitative measure of gender representation but lacks the nuance of qualitative analysis. Understanding the context could reveal more about the nature of gender bias, such as whether certain roles or characteristics are predominantly associated with a specific gender. Future work should incorporate contextual analysis to provide a richer and more comprehensive understanding of gender bias in text.
@@
Budget considerations: Using commercial, state-of-the-art LLMs in research can have a non-negligible financial cost. While we selected representative subsets from widely used benchmark datasets, a more extensive analysis could potentially provide more robust and generalizable findings at a cost.